# State Permutation Control in Non-Hermitian Multiqubit Systems with Suppressed Non-Adiabatic Transitions

### v1

@@ V Discussion
It is evident that by considering larger systems with $(2n+1)>3$ qubits, more complex permutation groups can be implemented in the system dynamics. Importantly, the ability to realize permutation groups in the time-dependent NHH in Eq. (1) enables dynamical state transfer or swapping for any arbitrary state, since any state can be expressed as a superposition of the NHH eigenstates. This is in striking contrast with previous studies on dynamical winding around or in the vicinity of EPs, where non-adiabatic transitions prevent bijective mappings over NHH eigenstates due to chiral state switching [21, 1].
@@
It is important to emphasize that when the initial state in the protocol described above is a superposition of the NHH eigenstates, the permuted eigenstates in that superposition accumulate an additional phase during state evolution. This phase generally comprises two components: the topological phase, arising from the encirclement of the EPs, and the dynamic phase. Such nontrivial phase accumulation can be further utilized in designing phase-dependent quantum protocols, akin to those employed in holonomic computation schemes [47].
@@
The NHH of the form in Eq. (1) with effective complex magnetic fields can be experimentally realized, e.g., through a Naimark dilation method, where the given NHH constitutes a part of a larger Hermitian system [50, 56, 57, 58, 59]. Moreover, because the pseudo-Hermiticity of the NHH $\hat{H}$ (Eq. (1)) does not undergo spontaneous symmetry breaking, the exact form of the dilated Hermitian Hamiltonian can be readily found [58] (see also Appendix B for details). This method can be experimentally implemented using platforms based on nitrogen-vacancy centers in diamond [57, 59] or superconducting quantum processors [60, 61].
Alternatively, the NHH can be simulated in open quantum optical systems with hard-core bosons, where a fermionic system is effectively mapped to a bosonic one [62, 63]. In that case, the complex magnetic field can act as a complex driving coherent field, which can be emulated using postselection techniques based on homodyne detection [51]. However, in contrast to the ‘deterministic’ dilation method, that approach is inherently probabilistic, since it relies on a post-selection of trajectories that requires continuous monitoring of the environment. Furthermore, the experimental simulation of the NHH can also be performed in classical and quantum photonic setups consisting of a network of coupled waveguides, whose coupling topology is identified with the Fock space of the NHH [1].
@@
The described protocol is robust against perturbations in the system that are sufficiently small to ensure that (i) the energy gap opening at the branch cut remains small enough that the state transfer is still feasible via Landau–Zener–Stückelberg–Majorana transitions [64], and (ii) the induced disorder does not trigger non-adiabatic transitions in the system dynamics. Indeed, in general, arbitrary disorder breaks the pseudo-Hermiticity of the NHH in Eq. (1), leading to a complex-valued spectrum and to system instabilities associated with this. We elaborate further on this in Appendix E.
@@
Furthermore, variations in the orbiting trajectory of the magnetic field in the $(x,y)$ plane do not affect the resulting state permutations, as long as the dynamical loop encloses the EPs. The protocol also remains stable when the modulated interaction terms $J_{mn}\approx 0$ (but not exactly zero) at $t=T/2$ in Eq. (IV.1), meaning that the different Riemann sheet pairs do not precisely intersect at that moment. This robustness is again attributed to the Landau–Zener–Stückelberg–Majorana transitions between the corresponding energy manifolds, ensuring that the observed state permutations are still achievable [42].
@@
We note that although in this work we specifically focused on the NHH in the form given in Eq. (1), other types of NHHs exhibiting various dynamic permutation groups in their eigenspace can also be identified in a similar manner. Our primary aim here was to highlight and describe the very possibility of the existence of such Hamiltonians that enable controlled and stable symmetric state switching in multiqubit systems via dynamical encirclement of EPs. Nevertheless, the formalism presented here can also be further extended to multiqubit systems with an even number of qubits. This could open up new possibilities for implementing genuine adiabatic quantum computing in non-Hermitian setups, e.g., realizing two-qubit quantum protocols in addition to the one- and three-qubit gates presented here.
@@ VI Conclusion
In this work, we have demonstrated that various quantum state permutations can be realized in multiqubit non-Hermitian systems by dynamically encircling EPs. This is accomplished by specifically selecting a subspace of the system parameters where the NHH exhibits a real-valued spectrum. Dynamically winding around EPs within this subspace suppresses non-adiabatic transitions in the system dynamics, thereby enabling precise and controlled state permutations.
As an illustrative example, we studied an NHH describing an interacting three-qubit system. We showed that appropriately time-modulating the system parameter space allows implementing a nontrivial permutation group acting on the eigenstates of the given NHH. Furthermore, the group generators enable symmetric switching between any two eigenstates of the NHH, regardless of the winding direction.
The NHHs analyzed can be experimentally realized across various platforms, including solid-state systems such as superconducting qubits or nitrogen-vacancy centers in diamond using the dilation method. They can also be simulated in open quantum bosonic systems subject to quadrature fields postselection, or emulated in classical photonic setups through networks of coupled cavities or waveguides.
The ability to perform controlled state permutations, while dynamically encircling EPs, introduces new opportunities for quantum state generation and transfer. In particular, the method described here may lead to novel experimental approaches to performing holonomic quantum operations that are based on non-Hermitian quantum physics. Successful realizations of such schemes are likely to require additional corrective controls to manage spurious transitions when performing operations on quasi-diabatic timescales. The current findings thus highlight the potential of non-Hermitian systems to advance quantum information and communication protocols, opening up novel avenues for state control and manipulation.
@@ Acknowledgements.
Here we provide further details on pseudo-Hermiticity of the NHH in Eq. (1) in the main text. Namely, we want to show that one can find such a Hermitian operator $\eta$ for which one always has
$\eta\hat{H}\eta^{-1}=\hat{H}^{\dagger}.$ (16)
@@
The states $|\nu_{i}\rangle$ are the left eigenvectors of the single qubit Hamiltonian, i.e., $H_{q}^{\dagger}|\nu_{i}\rangle\equiv|\nu_{i}\rangle$, with $H_{q}=B_{x}\sigma_{x}+B_{y}\sigma_{y}$, and where $B_{x,y}$ are given in Eq. (1). That is, the operator $\xi$ is a metric operator of the qubit Hamiltonian $H_{q}$. We recall that the left eigenvectors $\nu_{1,2}$ are complex conjugates of the right eigenvectors in Eq. (7).
@@
Because of this, the non-interacting part ($J=0$) of the NHH in Eq. (1), namely, $H_{\rm non}=\sum\limits_{k}^{N}\vec{{\boldsymbol{\sigma}}}^{k}\vec{\mathbb{B}}_{||}^{k}$, automatically fulfills the condition in Eq. (16), since the operator $\eta$ is a metric operator of this non-interacting term.
@@
On the other hand, the interacting term $H_{\rm int}=\sum J_{mn}\hat{\sigma}_{z}^{m}\hat{\sigma}_{z}^{n}$ is invariant under the action of the operator $\eta$ in Eq. (17), i.e., $\eta\hat{H}_{\rm int}\eta^{-1}=H_{\rm int}$, which can be checked by straightforward calculations.
As such, one attains
$\eta\hat{H}\eta^{-1}=\eta\hat{H}_{\rm non}\eta^{-1}+\eta\hat{H}_{\rm int}\eta^{-1}=\hat{H}_{\rm non}^{\dagger}+\hat{H}_{\rm int}.$ (18)
Because the $\hat{H}_{\rm int}$ is Hermitian by definition, one arrives at the pseudo-Hermitian condition in Eq. (16).
Appendix B Realizing the non-Hermitian Hamiltonian in Eq. (1) via the dilation method
@@
From the system of equations in Eq. (B) one attains a formal solution for the matrix $h_{1}$

### v2

@@ V Discussion
It is evident that by considering larger systems with $N>3$ qubits, more complex permutation groups can be implemented in the system dynamics. We note that even in the three-qubit system studied, other state permutations may be found by choosing different time-modulating functions of system parameters in Eq. (IV.1).
@@
Importantly, the ability to realize permutation groups in the time-dependent NHH in Eq. (1) enables dynamical state transfer or swapping for any arbitrary state, since any state can be expressed as a superposition of the NHH eigenstates. This is in striking contrast with previous studies on dynamical winding around or in the vicinity of EPs, where non-adiabatic transitions prevent bijective mappings over NHH eigenstates due to chiral state switching [21, 1].
@@
We would like to emphasize that when the initial state in the protocol described above is a superposition of the NHH eigenstates, the permuted eigenstates in that superposition accumulate an additional phase during state evolution. This phase generally comprises two components: the topological phase, arising from the encirclement of the EPs, and the dynamic phase. Such nontrivial phase accumulation [63, 64] can be further utilized in designing phase-dependent quantum protocols, akin to those employed in holonomic computation schemes [50, 51, 52].
We also reiterate that the divergent behavior of the spectrum near the EP, as indicated by Eq. (9), and associated with it infinite fields $\vec{\mathbb{B}}(x,y)$, does not affect the state-switching protocol proposed here. This is because the spectrum remains finite along the EP encircling trajectories. Moreover, as noted earlier in Section III.A, such divergences can be effectively circumvented by appropriately mapping the NHH onto a higher-dimensional NHH (see also Appendix B), thus allowing to access such EPs experimentally.
@@
The NHH of the form in Eq. (1) with effective complex-valued magnetic fields can be experimentally realized, e.g., through a Naimark dilation method, where the given NHH constitutes a part of a larger Hermitian system [55, 65, 66, 67, 68]. Moreover, because the pseudo-Hermiticity of the NHH $\hat{H}$ [(Eq. (1))] does not undergo spontaneous symmetry breaking, the exact form of the dilated Hermitian Hamiltonian can be readily found [67] (see Appendix D for details). This method can be experimentally implemented using platforms based on nitrogen-vacancy centers in diamond [66, 68] or superconducting quantum processors [69, 70].
Alternatively, the NHH can be simulated in open quantum optical systems operating in the hard-core boson regime, where a bosonic system can be effectively mapped to a spin system [71]. For example, in a lossy nonlinear cavity with strongly interacting photons, the system dynamics can be effectively mapped onto those of interacting qubits [72]. In that case, the complex magnetic field can act as a complex driving coherent field, which can be emulated using postselection techniques based on $\beta$-dyne detection [56] (see also Appendix E for details).
Furthermore, the experimental simulation of the NHH can also be performed in classical and quantum photonic setups consisting of real-space networks of dissipatively coupled waveguides, or even synthetic spaces, whose coupling topology is identified with the Fock space of the NHH [73, 74, 75, 76, 77].
@@
The described protocol is robust against perturbations in the system that are sufficiently small to ensure that (i) the energy gap opening at the branch cut remains small enough that the state transfer is still feasible via Landau–Zener–Stückelberg–Majorana transitions [78, 79], and (ii) the induced disorder does not trigger non-adiabatic transitions in the system dynamics. Indeed, in general, arbitrary disorder breaks the pseudo-Hermiticity of the NHH in Eq. (1), leading to a complex-valued spectrum and to system instabilities associated with this. We elaborate further on this in Appendix F.
@@
The protocol also remains stable when the modulated interaction terms $J_{kl}\approx 0$ (but not exactly zero) at $t=T/2$ in Eq. (IV.1), meaning that the different Riemann sheet pairs do not precisely intersect at that moment. This robustness is again attributed to the Landau–Zener–Stückelberg–Majorana transitions between the corresponding energy manifolds, ensuring that the observed state permutations are still achievable [44]. Furthermore, variations in the orbiting trajectory of the magnetic field in the $(x,y)$ plane do not affect the resulting state permutations, as long as the dynamical loop encloses the EPs and does not pass in its close proximity. However, we observe that for certain encircling loops and specific states, numerical simulations may exhibit stiffness. We discuss this difficulty and the way to overcome it in more detail in Appendix G.
@@
We note that although in this work we specifically focused on the NHH in the form given in Eq. (1), other types of NHHs exhibiting various dynamic permutation groups in their eigenspace can also be identified in a similar manner. Our primary aim here was to highlight and describe the very possibility of the existence of such Hamiltonians that enable controlled and stable symmetric state switching in multiqubit systems via dynamical encirclement of EPs. However, the formalism presented here can also be further extended to multiqubit systems with any number of qubits. This could open up new possibilities for implementing genuine adiabatic quantum computing in non-Hermitian setups, e.g., realizing two-qubit quantum protocols in addition to the one- and three-qubit gates presented here.
@@ VI Conclusion
In this work, we have demonstrated that genuine topological quantum state permutations can indeed be realized in multiqubit non-Hermitian systems by dynamically encircling EPs, by converting the Hamiltonian to a pseudo-Hermitian Hamiltonian. This is accomplished by adding complex fields to the system and specifically selecting a subspace where the NHH exhibits a real-valued spectrum and hosts a novel type of EPs defined on hyperboloid manifolds in the system parameter space. We term these exception points ’twisted EPs’. Dynamically winding around these EPs allows the suppression of non-adiabatic transitions in the system dynamics, thereby enabling precise and controlled quantum state permutations.
As an illustrative example, in this work we studied an NHH describing an interacting three-qubit system. We showed that appropriately time-modulating the system parameter space allows the implementation of a nontrivial non-Abelian permutation group acting on the eigenstates of the given NHH. Furthermore, the group generators enable symmetric switching between any two eigenstates of the NHH, regardless of the winding direction, removing the mode chirality imposed by non-adiabatic transitions in conventional non-Hermitian systems.
The NHHs with twisted EPs that are proposed and analyzed in this work can be experimentally realized across various platforms, including solid-state systems, such as superconducting qubits or nitrogen-vacancy centers in diamond using the dilation method [69, 70, 66, 68]. They can also be simulated in the monitored open quantum interacting bosonic systems by performing a photon postselection procedure, or emulated in classical photonic setups through networks of coupled cavities or waveguides [56, 73, 74, 75, 76, 77].
The ability to perform controlled state permutations, while dynamically encircling EPs, introduces new opportunities for quantum state generation and transfer. In particular, the method described here may lead to novel experimental approaches to performing holonomic quantum operations that are based on non-Hermitian quantum physics. Indeed, current non-Hermitian holonomic protocols are severely limited by non-adiabatic transitions between energy eigenstates and the losses associated with these [80]. Successful realizations of such approaches are likely to require additional corrective controls to manage spurious transitions when performing operations on quasi-diabatic timescales. In constrast, the present approach exhibits no non-adiabatic losses and can be implemented by standard control methods. The current findings thus highlight the potential of non-Hermitian systems to advance quantum information and communication protocols, opening up novel avenues for state control and manipulation.
@@ Acknowledgements.
Here we provide further details on the pseudo-Hermiticity of the NHH in Eq. (1) in the main text. Specifically, we want to show that one can find a Hermitian operator $\eta$ for which one always has
$\eta\hat{H}\eta^{-1}=\hat{H}^{\dagger},$ (20)
@@
where the states $|\nu_{i}\rangle$ are the left eigenvectors of the single qubit Hamiltonian $H_{q}=B_{x}\sigma_{x}+B_{y}\sigma_{y}$, with $B_{x,y}$ given by Eq. (1) - Eq. (3), i.e., $H_{q}^{\dagger}|\nu_{i}\rangle\equiv|\nu_{i}\rangle$. That is, the operator $\xi$ is a metric operator of the qubit Hamiltonian $H_{q}$. We recall that the left eigenvectors $\nu_{1,2}$ are complex conjugates of the right eigenvectors in Eq. (8).
@@
Because of this property, the non-interacting part ($\textbf{J}=0$) of the NHH in Eq. (1), namely, $H_{\rm non}=\sum\limits_{k}^{N}\vec{{\boldsymbol{\sigma}}}^{k}\vec{\mathbb{B}}_{||}^{k}$, automatically fulfills the condition in Eq. (20), since the operator $\eta$ is a metric operator of this non-interacting term.
@@
On the other hand, the interacting term $H_{\rm int}=\sum J_{kl}\hat{\sigma}_{z}^{k}\hat{\sigma}_{z}^{l}$ is invariant under the action of the operator $\eta$ in Eq. (21), i.e., $\eta\hat{H}_{\rm int}\eta^{-1}=H_{\rm int}$, which can be checked by straightforward calculations.
Consequently, we obtain
$\eta\hat{H}\eta^{-1}=\eta\hat{H}_{\rm non}\eta^{-1}+\eta\hat{H}_{\rm int}\eta^{-1}=\hat{H}_{\rm non}^{\dagger}+\hat{H}_{\rm int}.$
Because $\hat{H}_{\rm int}$ is Hermitian by definition, one then arrives at the pseudo-Hermitian condition in Eq. (20).
Appendix B Resolving the problem of exceptional points at infinity for single qubits in Eq. (9)
Here, we briefly address the apparent issue of spectral divergence at EPs in the non-interacting qubits NHH, according to Eq. (9). This problem can be resolved using the framework of complex projective spaces [59], which provides an elegant method to handle infinities. In particular, what appears as an infinity in a given space may correspond to a well-defined point in a higher-dimensional space.
Such a higher-dimensional $3\times 3$ NHH can be introduced as follows
where $\gamma=\Delta\sin\left({\rm Re}[\phi]\right)$. The first two diagonal elements of the NHH are continuous functions on the plane $(x,y)$. The third element is also a continuous function except the interval $x=0,|y|\leq 1]$, where it becomes rather a piecewise function, i.e., it is not differentiable but remains finite on that interval.
$\lambda_{1}^{\rm EP}=\lambda_{2}^{\rm EP}=0.$ (25)
Figure 6: This Figure is an extension of Figs. 4(d) and 5(d), i.e., the calculated state transfer fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between evolving state $|\psi(t)\rangle$ and one of the eight initial eigenstates $|\psi_{k}\rangle$ of the three-qubit NHH. The panels in the plot are related to different combinations of the modulated system parameters $J_{kl}$ in Eq. (IV.1), and which thus correspond to permutation generators in Eq. (IV.3). The panels (a),(b),(c) and (d) correspond to the permutation generators $p_{3}$, $p_{4}$, $p_{5}$ and $p_{6}$, respectively. The semi-minor axis and semi-major axis of the encircling ellipse are $r_{x}=1$ and $r_{y}=2$, respectively. The remaining system parameters are the same as in Figs. 4(d) and 5(d).
Figure 7: Schematic representation of implementing an element $p_{3}$ of the permutation group $\cal G$ in Eq. (IV.3) in the eigenspace of a three-qubit system governed by the NHH in Eq. (1), while dynamically spiraling EPs over a winding period $T$. The system parameters are modulated according to Eq. (IV.1). This figure also corresponds to the fidelity graph shown in Fig. 6(a). (a) Initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold [same as in Figs. 4(a) and 5(a)]. (b) The position of evolving eigenstates at time $0<t<T/2$. (c) Zoomed in incaption from panel (b), showing the evolution of the eigenstates $\psi_{4}$ (red sphere), $\psi_{5}$ (cyan sphere) and $\psi_{6}$ (magenta sphere), which pass through various Riemann energy manifolds. (d,e,f) The evolution of the eigenstates at times $t=T/2$, $T/2<t<T$, and $t=T$, respectively. See the main text in Appendix C.1 for more details.
Figure 8: Schematic representation of implementing an element $p_{4}$ of the permutation group $\cal G$ in Eq. (IV.3) in the eigenspace of a three-qubit system governed by the NHH in Eq. (1), while dynamically spiraling EPs over a winding period $T$. The system parameters are modulated according to Eq. (IV.1). This figure also corresponds to the fidelity graph shown in Fig. 6(b). (a) Initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold [same as in Figs. 4(a) and 5(a)]. (b) The position of evolving eigenstates at time $0<t<T/2$. (c) Zoomed in incaption from panel (b), showing the evolution of the eigenstates $\psi_{2}$ (green sphere) and $\psi_{3}$ (yellow sphere), whose trajectories along the corresponding Riemann manifolds are exchanged at the touching point of these energy sheets. (d,e,f) The evolution of the eigenstates at times $t=T/2$, $T/2<t<T$, and $t=T$, respectively. See the main text in Appendix C.2 for more details.
Figure 9: Schematic representation of implementing an element $p_{5}$ of the permutation group $\cal G$ in Eq. (IV.3) in the eigenspace of a three-qubit system governed by the NHH in Eq. (1), while dynamically spiraling EPs over a winding period $T$. The system parameters are modulated according to Eq. (IV.1). This figure also corresponds to the fidelity graph shown in Fig. 6(c). (a) Initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold [same as in Figs. 4(a) and 5(a)]. (b,c,d) The position of the evolving eigenstates at times $0<t<T/2$, $t=T/2$, and $t=T$, respectively. See the main text in Appendix C.3 for more details.
Figure 10: Schematic representation of implementing an element $p_{6}$ of the permutation group $\cal G$ in Eq. (IV.3) in the eigenspace of a three-qubit system governed by the NHH in Eq. (1), while dynamically spiraling EPs over a winding period $T$. The system parameters are modulated according to Eq. (IV.1). This figure also corresponds to the fidelity graph shown in Fig. 6(d). (a) Initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold (same as in Figs. 4 and 5). (b,c,d) The position of the evolving eigenstates at times $0<t<T/2$, $t=T/2$, and $t=T$, respectively. See the main text in Appendix C.4 for more details.
Appendix C Calculated fidelity for other state permutations according to Eq. (IV.3)
In Figs. 4(d) and 5(d) we plotted the fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between the evolving state $|\psi(t)\rangle$ and one of the eight initial eigenstates $|\psi_{k}\rangle$ of the three-qubit NHH that correspond to the first generators of the permutation group $\cal G$ in Eq. (IV.3). For completeness, here in Fig. 6 we plot the fidelity of the remaining permutation generators, thus confirming their form in Eq. (IV.3). Visualizations of he state dynamics corresponding to these permutations are shown in Figs. 7 - 10 and the various eigenstate flips produced by the permutation generators $p_{i}$ are summarized in Table 1. From this Table it is clear that one can always find such a combination of modulated system parameters in Eq. (IV.1) that enables a state transfer between any two given eigenstates in one or more full cycles $T$.
Below we analyze in detail the evolution of eigenstates corresponding to each panel in Fig. 6.
C.1 Visualization of the eigenstates evolution corresponding to the element $p_{3}$ of the permutation group $\cal G$ in Eq. (IV.3)
We begin with the state dynamics illustrated in Fig. 7, whose fidelity evolution is shown in Fig. 6(a), corresponding to the permutation group element $p_{3}$ of $\cal G$ defined in Eq. (IV.3).
At $t=0$, the initial eigenstates $\psi_{i}$ occupy their respective Riemann eigenenergy sheets $E_{i}$ in the Hermitian phase of the non-Hermitian Hamiltonian $\hat{H}$, where the complex magnetic field $\boldsymbol{\vec{B}}=0$, i.e., $(x=0,y=0)$ in Fig. 7(a). The eigenstates $\psi_{i}$ are represented as colored spheres, consistent with Figs. 4 and 5: $\psi_{1}$ (blue), $\psi_{2}$ (green), $\psi_{3}$ (yellow), $\psi_{4}$ (red), $\psi_{5}$ (cyan), $\psi_{6}$ (magenta), $\psi_{7}$ (black), and $\psi_{8}$ (brown).
As the system parameters are modulated according to Eq. (IV.1), including the time-dependent couplings $J_{13}(t)$ and $J_{23}(t)$, the Riemann sheets begin to deform, showing increasing bending, and also shift along the $z$-axis, as shown in Fig. 7(b). At an intermediate time $0<t<T/2$ within the range [1<x<0, 0<y<2][-1<x<0,\ 0<y<2], the energy manifolds $E_{4}$, $E_{5}$, and $E_{6}$ undergo mutual permutation, according to the adopted convention in Eq. (7), [see Fig. 7(c), which is a zoom-in incaption of panel (b)]. Consequently, the associated eigenstates are reshuffled: $\psi_{4}$ begins evolving along $E_{5}$, $\psi_{5}$ follows $E_{6}$, and $\psi_{6}$ follows $E_{4}$.
At the midpoint of the evolution, $t=T/2$ ($x=0$, $y=2$), several Riemann sheets become coupled: $E_{1}\leftrightarrow E_{3}$, $E_{2}\leftrightarrow E_{4}$, $E_{5}\leftrightarrow E_{8}$, and $E_{6}\leftrightarrow E_{7}$ [Fig. 7(d)]. These couplings cause the eigenstates to switch to different sheets: $\psi_{1}\rightarrow E_{3}$, $\psi_{2}\rightarrow E_{4}$, $\psi_{3}\rightarrow E_{1}$, $\psi_{4}$ (on $E_{5}$) $\rightarrow E_{8}$, $\psi_{5}$ (on $E_{6}$) $\rightarrow E_{7}$, $\psi_{6}$ (on $E_{4}$) $\rightarrow E_{2}$, $\psi_{7}\rightarrow E_{6}$, and $\psi_{8}\rightarrow E_{5}$.
Subsequently, during the second half of the cycle, for $T/2<t<T$, shown in Fig. 7(e) an identical permutation among $E_{4}$, $E_{5}$, and $E_{6}$ occurs, mirroring the earlier rearrangement in Fig. 7(c). This time, the permutation involves the states $\psi_{2}$, $\psi_{7}$, and $\psi_{8}$.
As a result of the complete encirclement around the EPs, as depicted in Fig. 7(f), one realizes the permutation element:
$p_{3}=(\psi_{1},\psi_{5})(\psi_{2},\psi_{6})(\psi_{3},\psi_{7})(\psi_{4},\psi_{8}),$ (26)
according to Eq. (IV.3).
C.2 Visualization of the eigenstates evolution corresponding to the element $p_{4}$ of the permutation group $\cal G$ in Eq. (IV.3)
By modulating the magnetic field $\boldsymbol{\vec{B}}(x,y)$ along with the time-dependent qubit coupling $J_{12}(t)$, according to Eq. (IV.1), one can realize the eigenstate permutation associated with the group element $p_{4}$ of the permutation group $\cal G$ defined in Eq. (IV.3). The resulting evolution of the eigenstates corresponding to $p_{4}$ is illustrated in Fig. 8.
As in Fig. 7(a), the initial eigenstates at $t=0$ are aligned with their respective Riemann energy sheets in the Hermitian phase $\boldsymbol{\vec{B}}(x,y)=0$, as shown in Fig. 8(a). At an intermediate stage, $0<t<T/2$, in the domain [1<x<0, 0<y<2][-1<x<0,\ 0<y<2], the deformation of the energy sheets $E_{2}$ and $E_{3}$ due to the modulation of $J_{12}(t)$ may cause these surfaces to come into proximity, where they touch each other [see Fig. 8(c), which is a zoom-in of the rectangular region highlighted in panel (b)]. As a result, the corresponding eigenstates undergo a permutation: $\psi_{2}$ begins evolving along $E_{3}$, while $\psi_{3}$ transitions to $E_{2}$.
At the halfway point of the parameter cycle, $t=T/2$ (x=0,y=2x=0,\ y=2), several Riemann sheets become coupled, though with a different pairing compared to the previous case. The coupled energy surfaces are: $E_{1}\leftrightarrow E_{2}$, $E_{3}\leftrightarrow E_{6}$, $E_{4}\leftrightarrow E_{5}$, and $E_{7}\leftrightarrow E_{8}$ [Fig. 8(d)]. These couplings induce the following transitions among the eigenstates: $\psi_{1}\rightarrow E_{2}$, $\psi_{2}$ (on $E_{3}$) $\rightarrow E_{6}$, $\psi_{3}$ (on $E_{2}$) $\rightarrow E_{1}$, $\psi_{4}\rightarrow E_{5}$, $\psi_{5}\rightarrow E_{4}$, $\psi_{6}\rightarrow E_{3}$, $\psi_{7}\rightarrow E_{8}$, and $\psi_{8}\rightarrow E_{7}$.
In the second half of the encirclement, for $T/2<t<T$, a similar exchange between $E_{2}$ and $E_{3}$ takes place, mirroring the earlier scenario in Fig. 8(c). This leads to an additional permutation between $\psi_{1}$ and $\psi_{6}$.
At the end of the full cycle ($t=T$), the final configuration of the eigenstates corresponds to the permutation element:
$p_{4}=(\psi_{1},\psi_{3})(\psi_{2},\psi_{6})(\psi_{4},\psi_{5})(\psi_{7},\psi_{8}),$ (27)
when compared to the initial state ordering.
C.3 Visualization of the eigenstates evolution corresponding to the element $p_{5}$ of the permutation group $\cal G$ in Eq. (IV.3)
The implementation of the eigenstate permutation $p_{5}$ from the group $\cal G$ requires, in addition to modulating the magnetic field $\boldsymbol{\vec{B}}(x,y)$, the simultaneous modulation of either the qubit couplings $J_{12}(t)$ and $J_{13}(t)$, or the pair $J_{12}(t)$ and $J_{23}(t)$, in accordance with Eqs. (IV.1) and (IV.3). We illustrate the corresponding state evolution in Fig. 9.
In contrast to the permutations $p_{3}$ and $p_{4}$, no eigenstate transitions occur during the intermediate intervals $0<t<T/2$ or $T/2<t<T$ [see also Fig. 9(b)]. Instead, the state evolution is entirely governed by the couplings at the midpoint of the winding cycle, $t=T/2$. At this point, the following Riemann sheet couplings are observed: $E_{1}\leftrightarrow E_{3}$, $E_{2}\leftrightarrow E_{4}$, $E_{5}\leftrightarrow E_{7}$, and $E_{6}\leftrightarrow E_{8}$ [see Fig. 9(c)]. These couplings induce the corresponding eigenstate transitions: $\psi_{1}\rightarrow E_{3}$, $\psi_{2}\rightarrow E_{4}$, $\psi_{3}\rightarrow E_{1}$, $\psi_{4}\rightarrow E_{2}$, $\psi_{5}\rightarrow E_{7}$, $\psi_{6}\rightarrow E_{8}$, $\psi_{7}\rightarrow E_{5}$, and $\psi_{8}\rightarrow E_{6}$. This new configuration of eigenstates remains unchanged for the rest of the evolution until the end of the winding period $t=T$, thus leading to the permutation element $p_{5}$ defined in Eq. (IV.3):
$p_{5}=(\psi_{1},\psi_{3})(\psi_{2},\psi_{4})(\psi_{5},\psi_{7})(\psi_{6},\psi_{8}).$ (28)
C.4 Visualization of the eigenstates evolution corresponding to the element $p_{6}$ of the permutation group $\cal G$ in Eq. (IV.3)
To realize the eigenstate permutation $p_{6}$, one must simultaneously modulate the magnetic field $\boldsymbol{\vec{B}}(x,y)$ and all three qubit couplings $J_{12}(t)$, $J_{13}(t)$, and $J_{23}(t)$, according to Eqs. (IV.1) and (IV.3). The corresponding eigenstate dynamics are shown in Fig. 10.
As in the case of the permutation element $p_{5}$, the final configuration of eigenstates is solely determined by the couplings between Riemann sheets at the midpoint of the evolution, $t=T/2$, induced by the modulating qubit couplings. Specifically, Fig. 10(c) shows that the following energy surfaces become coupled and degenerate: $E_{1}\leftrightarrow E_{8}$, $E_{2}\leftrightarrow E_{6}$, $E_{3}\leftrightarrow E_{7}$, and $E_{4}\leftrightarrow E_{5}$. These couplings result in the corresponding transitions between eigenstates: $\psi_{1}\rightarrow E_{8}$, $\psi_{2}\rightarrow E_{6}$, $\psi_{3}\rightarrow E_{7}$, $\psi_{4}\rightarrow E_{5}$, $\psi_{5}\rightarrow E_{4}$, $\psi_{6}\rightarrow E_{2}$, $\psi_{7}\rightarrow E_{3}$, and $\psi_{8}\rightarrow E_{1}$. This evolution leads to the realization of the permutation element $p_{6}$ of the group $\cal G$, given by:
$p_{6}=(\psi_{1},\psi_{8})(\psi_{2},\psi_{6})(\psi_{3},\psi_{7})(\psi_{4},\psi_{5}).$ (29)
$\psi_{1}$ $p_{2}$ $p_{3},p_{4},p_{5}$ $p_{1}$ $p_{6}$
$\psi_{2}$ $p_{2}$ $p_{5}$ $p_{1},p_{3},p_{4},p_{6}$
$\psi_{3}$ $p_{3},p_{4},p_{5}$ $p_{1}$ $p_{2}$ $p_{6}$
$\psi_{4}$ $p_{5}$ $p_{1}$ $p_{2},p_{4}$ $p_{3}$
$\psi_{5}$ $p_{1}$ $p_{2},p_{4}$ $p_{3},p_{5}$
$\psi_{6}$ $p_{1},p_{3},p_{4},p_{6}$ $p_{2}$ $p_{5}$
$\psi_{7}$ $p_{6}$ $p_{3},p_{5}$ $p_{1},p_{2},p_{4}$
$\psi_{8}$ $p_{6}$ $p_{3}$ $p_{5}$ $p_{1},p_{2},p_{4}$
Appendix D Realizing the non-Hermitian Hamiltonian in Eq. (1) via the dilation method
@@
From the system of equations in Eq. (D) one attains a formal solution for the matrix $h_{1}$